what is the equation of the function which has a graph that is the reflection across the x-axis of the graph of f(x)=x^2+2x-5 ?
Answer by ikleyn(52797)

.
what is the equation of the function which has a graph that is the reflection across the x-axis of the graph of f(x) = x^2+2x-5 ?
~~~~~~~~~~~~~~~~~~~

The function for the reflected graph is


g(x) = - f(x):   you only need to change the sign at the function to the opposite one.


So,  g(x) = -(x^2+2x-5) = -x^2 - 2x + 5.


It is how the reflection about the x-axis works: (x,y) ----> (x,-y).